OpenStudy (anonymous):

A man walks 5 kilometers in an hour. How many seconds will it take him to walk 150 yards?

OpenStudy (anonymous):

150 yards = 137.16 meter and 5 km/hr = 25/18 m/sec time = distance / speed time = 137.16*18/25 = 98.75 sec